OpenStudy (epicteatime):

2/3*(x - 2) = 4x pls tell me how to do!!

OpenStudy (yttrium):

Is it \[\frac{ 2 }{ 3 (x-2) } = 4x\]

OpenStudy (epicteatime):

no um its \[2/3 \times (x-2) = 4x \]

OpenStudy (epicteatime):

2 and 3 is a fraction

OpenStudy (anonymous):

given that 2/3*(x - 2) = 4x then we can write it as 2(x-2)=3*4x 2x-4=12x -4=12x-2x -4=10x then x=-4/10

OpenStudy (anonymous):

did you get it.....@Epicteatime

OpenStudy (texaschic101):

2/3(x - 2) = 4x -- distribute through the parenthesis 2/3x - 4/3 = 4x -- multiply by the LCD, which is 3 2x - 4 = 12x -- subtract 2x from both sides -4 = 12x - 2x -- combine like terms -4 = 10x -- divide by 10 -4/10 = x -- reduce - 2/5 = x check... 2/3(x - 2) = 4x 2/3(-2/5 - 2) = 4(-2/5) 2/3(-2/5 - 10/5) = - 8/5 2/3(-12/5) = -8/5 - 24/15 = -8/5 -8/5 = -8/5 (correct) x = -2/5

OpenStudy (epicteatime):

YUP THANX GUYS!!